OpenStudy (chihiroasleaf):

[Partial Differential Equation] Show that each of the following equations has a solution of the form \( u(x,y)=f(ax+by) \) for a proper choice of constant \( a,b\) . (a) \(u_{x}+3u_{y}=0\) (b) \(3u_{x}−7u_{y}=0\) (c)\( 2u_{x}+πu_{y}=0\)

OpenStudy (anonymous):

ok this is a matter of substitution

OpenStudy (anonymous):

solving for part (a)\[a\frac{ \delta }{ \delta x }(ax+by)+3b\frac{ \delta }{ \delta y }(ax+by)=0\]

OpenStudy (anonymous):

or\[af \prime + 3b f \prime =0\]

OpenStudy (anonymous):

\[\ \therefore a=-3b\]

OpenStudy (chihiroasleaf):

here.., we conclude \(a = -3b\) since by taking this value, the equation will become true and the function \(f(ax+by)\) is the solution of the equation, right?

OpenStudy (anonymous):

and that satisfies the solution, so you try part b, and c, using the same method applied above

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i hope this helps

OpenStudy (chihiroasleaf):

part (b) \(3a-7b =0\) (c) \(2a+\pi b=0 \) the reason will be the same as part (a), right?

OpenStudy (anonymous):

yes

OpenStudy (chihiroasleaf):

ok..., thanks a lot... :)

OpenStudy (anonymous):

you're welcome